Betty is three times as old as her son now. In 12 years, Betty's age will be three years less than twice her son's age. How old is Betty’s son now?
Answer by ikleyn(52776)

.

Let x be the son's age now.


Then Betty's age is 3x years, according to the condition.


In 12 years


    3x + 12 = 2*(x+12) - 3


Simplify and solve


    3x + 12 = 2x + 24 - 3

    3x - 2x = 24 - 3 - 12

     x      = 9.


ANSWER.  The son is 9 years old now.

Solved.

--------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.